Suppose we estimate a regression of the gallons of paint sold on the price of each gallon and find the following estimated regression line:
y = 100 - 10x.
Predict how many gallons would be sold if the price is $5 per gallon?
Answer #1

Price is 5 implies x= 5

Put x= 5 into y= 100-10x

y= 100-10(5) = 50

Estimated sales are 50 at Price of 5.